# Evaluation for Omni MLLMs

### v1

@@ 1 Introduction
Multimodal large language models (MLLMs) have recently made impressive progress in bridging vision, language, and audio (Yin et al., 2024; Song et al., 2025; Cheng et al., 2025). While early benchmarks primarily focused on image-text alignment or visual reasoning (Xu et al., 2025a; Chen et al., 2024a; Yue et al., 2024a), the integration of video and audio presents a quite different challenge: models must jointly process long temporal sequences, dynamic scene transitions, and complementary acoustic cues. Despite rapid advances, evaluation of MLLMs on audio-visual reasoning remains underdeveloped. Existing benchmarks (Li et al., 2024a; Hong et al., 2025) often (i) focus on short video clips that underrepresent long-term temporal dependencies, (ii) emphasize a single modality (e.g., vision) while treating audio as auxiliary or optional. As a result, current evaluations fail to capture the challenges inherent to comprehensive video understanding, where audio and vision must be integrated consistently and logically to support robust inference.
@@ 2.5 Dataset Statistics
OmniBench (Li et al., 2024b) I+A 56.1 R Sp+So+Mu / MC
Daily-Omni (Zhou et al., 2025) V+A 47.5 R Sp+So+Mu 30\60 MC
@@
Table 2: Comparisons between different benchmarks. V, I, A for modality represent video, image and audio. Qwen2.5-Omni represents the performance of Qwen2.5-Omni-7B on these benchmarks. Multiple Domains signifies whether the video includes diverse domains. R and S in Video Type denote real-world and synthetic data. Sp, So, and Mu represent Speech, Sound, and Music for Audio Type, respectively. Video Duration represents the duration in seconds. MC, CLS for Answer Type indicate Multiple Choice and Classification from fixed vocabulary, respectively.
@@ 2.6 Dataset Comparison
In Table 2, we compare OmniVideoBench with existing benchmarks. While AV-Odyssey and OmniBench handle only single images, OmniVideoBench targets more challenging videos, ranging from a few seconds to 30 minutes, far exceeding the scope of other benchmarks. Besides, in Figure 4, OmniVideoBench imposes stringent demands on models’ cross-modal reasoning capabilities. Even the Qwen2.5-VL series models, which achieve 40.68% accuracy in the Daily-Omni (Zhou et al., 2025), perform near random levels on OmniVideoBench. This demonstrates that relying solely on pure visual information is insufficient. Moreover, OmniVideoBench presents a significantly greater challenge than existing omni-modal benchmarks. The widely used Qwen2.5-Omni-7B performs close to random guessing, highlighting the challenges of OmniVideoBench.
@@ 3.1 Baseline Models
We evaluate open-source MLLMs (i.e., Qwen3-Omni series (Xu et al., 2025b), Qwen2.5-Omni series (Xu et al., 2025c), Baichuan-Omni-1.5 (Li et al., 2025), HumanOmni (Zhao et al., 2025), MiniCPM-o (Yao et al., 2024), VideoLLaMA2 (Cheng et al., 2024)), and various closed-source MLLMs (i.e., Gemini-2.5-Pro, Gemini-2.5-Flash (Comanici et al., 2025), and Gemini-2.0-Flash). We also evaluate the Qwen2.5-VL series (Bai et al., 2025) and DeepSeek-V3.1 (Liu et al., 2024a).
Models Audio Type Video Duration Avg. Music Sound Speech (0,1] min (1,5] min (5,10] min (10,30] min Omni-Modal Language Models (With Visual and Audio) Gemini-2.5-Pro 38.46 57.72 61.66 57.83 64.43 55.02 55.94 58.90 Gemini-2.5-Flash 39.56 57.04 53.17 55.42 55.10 47.37 52.11 52.40 Gemini-2.0-Flash 29.67 40.27 43.21 49.40 43.15 41.05 34.87 41.50 Qwen3-Omni-30B-A3B 37.36 34.67 39.26 45.78 37.03 38.86 35.11 38.40 Baichuan-Omni-1.5 24.18 31.33 31.36 28.92 31.78 28.38 32.44 30.70 HumanOmni-7B 20.87 31.08 31.61 36.57 29.36 29.60 29.25 30.50 MiniCPM-o 27.47 28.57 30.24 31.43 28.49 34.53 26.15 29.70 Qwen2.5-Omni-7B 23.07 25.33 30.70 41.57 27.41 25.33 26.72 29.30 VideoLLaMA2-7B 26.37 30.67 29.25 32.00 28.20 29.60 28.29 29.20 Omni-Modal Language Models (Visual Only) Gemini-2.0-Flash 25.27 36.67 30.99 33.73 35.86 32.75 22.48 31.30 Qwen2.5-Omni-7B 27.47 26.67 26.22 28.31 27.11 24.45 25.95 26.40 Visual Language Models (Visual Only) Qwen2.5-VL-32B 32.97 32.00 31.49 38.55 31.20 29.26 30.53 31.80 Qwen2.5-VL-7B 29.67 31.33 29.51 25.90 30.03 31.88 30.15 29.80 Qwen2.5-VL-72B 26.37 29.33 29.91 33.13 30.03 31.88 24.43 29.50 Baseline LLMs DeepSeek-V3.1 28.57 26.17 27.28 30.91 27.57 25.00 26.44 27.60
@@ 3.3 Further Analysis
Models Open-ended QA MCQ
Gemini-2.0-Flash 27.06 41.50
Qwen2.5-Omni-7B 17.25 29.30
Effect of Different Numbers of Frames.
@@ 4 Related Works
Omni-Understanding MLLMs. The development of MLLMs (Chen et al., 2022; Awadalla et al., 2023; Liu et al., 2023) began with a foundational focus on integrating the two primary modalities of vision and language. A recent paradigm shift aims to develop Omni-modal MLLMs capable of processing and generating information across an arbitrary combination of modalities (“Any-to-Any"). This approach positions the LLM as a central cognitive engine, unifying diverse data types like audio, video, and text within its semantic space (Liu et al., 2024b; Yuan et al., 2025). This has driven a move from integrating pre-trained unimodal components towards developing “natively multimodal" architectures trained from the ground up, as exemplified by models like GPT-4o (Hurst et al., 2024). This ambition is showcased by state-of-the-art models (Xu et al., 2025c; Zhao et al., 2025; Li et al., 2024c; 2025; Yao et al., 2024; Sun et al., 2025; Liu et al., 2025b), which pioneer end-to-end streaming capabilities for simultaneously processing video and audio to generate text and speech. At the forefront of this paradigm, proprietary models like Gemini series (Team, 2024; Comanici et al., 2025) demonstrate pinnacle performance, powered by a natively multimodal design and a massive context window that together unlock superior understanding of complex, interwoven data streams.
MLLM Benchmarks. The landscape of MLLM evaluation has matured significantly, evolving from foundational perception benchmarks (Liu et al., 2024c; Li et al., 2024a; Yu et al., 2024a; b; Chen et al., 2024b; Jiang et al., 2025) to more sophisticated frameworks (He et al., 2025; Du et al., 2025). Recent efforts probe deeper cognitive abilities, with MLLM-Bench (Ge et al., 2025) assessing a hierarchy of cognitive skills. MMMU (Yue et al., 2023) and MMMU-Pro (Yue et al., 2024b) challenging models with expert-level, multi-disciplinary reasoning under stricter protocols like vision-only inputs. Simultaneously, evaluation has specialized into high-stakes domains such as finance (Gan et al., 2024) and medicine (Chen et al., 2024c). For video, some benchmarks (Wang et al., 2019; Li et al., 2021; 2023; Fang et al., 2024; Wu et al., 2024) now focus on the critical challenge of long-context temporal understanding (Liu et al., 2025c), revealing key limitations in current models.
@@ 6 Contributions
  • Jiafu Tang, Nanjing University

### v2

@@ 1 Introduction
Multimodal large language models (MLLMs) have recently made impressive progress in bridging vision, language, and audio (Yin et al., 2024; Song et al., 2025; Cheng et al., 2025). While early benchmarks primarily focused on image-text alignment or visual reasoning (Xu et al., 2025c; Chen et al., 2024c; Yue et al., 2024a), the integration of video and audio presents a quite different challenge: models must jointly process long temporal sequences, dynamic scene transitions, and complementary acoustic cues. Despite rapid advances, evaluation of MLLMs on audio-visual reasoning remains underdeveloped. Existing benchmarks (Li et al., 2024a; Hong et al., 2025) often (i) focus on short video clips that underrepresent long-term temporal dependencies, (ii) emphasize a single modality (e.g., vision) while treating audio as auxiliary or optional. As a result, current evaluations fail to capture the challenges inherent to comprehensive video understanding, where audio and vision must be integrated consistently and logically to support robust inference.
@@ 2.5 Dataset Statistics
AVTRUSTBENCH (Chowdhury et al., 2025) V+A / R+S Sp+So+Mu 10\60 MC
MMAU (Sakshi et al., 2024) A 71.0 / Sp+So+Mu / MC
DAVE (Radevski et al., 2025) V+A 31.0 R+S So $\leq 60$ MC
@@
AVHBench (Judgement) (Sung-Bin et al., 2024) V+A 74.7 R+S So 10 CLS
OmniBench (Li et al., 2024c) I+A 56.1 R Sp+So+Mu / MC
Daily-Omni (Zhou et al., 2025) V+A 47.5 R Sp+So+Mu 30–60 MC
@@
Table 2: Comparisons between different benchmarks and datasets. V, I, A for modality represent video, image and audio. Qwen2.5-Omni represents the performance of Qwen2.5-Omni-7B on these benchmarks. Multiple Domains signifies whether the video includes diverse domains. R and S in Video Type denote real-world and synthetic data. Sp, So, and Mu represent Speech, Sound, and Music for Audio Type, respectively. Video Duration represents the duration in seconds. MC, CLS for Answer Type indicate Multiple Choice and Classification from fixed vocabulary, respectively.
@@ 2.6 Dataset Comparison
In Table 2, we compare OmniVideoBench with representative audio-video benchmarks. While AV-Odyssey (Gong et al., 2024) and OmniBench (Li et al., 2024c) operate on single images, OmniVideoBench targets substantially more challenging videos with durations ranging from a few seconds to 30 minutes. Recent benchmarks such as AVTrustBench (Chowdhury et al., 2025), DAVE (Radevski et al., 2025), and MMAU (Sakshi et al., 2024) begin emphasizing audio–video coordination, but typically focus on specific capabilities or short clips. AVHBench (Sung-Bin et al., 2024) also evaluates audiovisual consistency, yet its tasks remain centered on shorter videos and hallucination detection. OmniVideoBench, by contrast, expands the scope to diverse video types, broader temporal spans, and fine-grained cross-modal reasoning, capturing richer dependencies between audio and vision. Compared to Daily-Omni (Zhou et al., 2025) and WorldSense (Hong et al., 2025), which also utilize multi-domain videos, OmniVideoBench places greater emphasis on explicit audiovisual collaboration. For instance, disabling audio causes Gemini-2.0-Flash’s performance to plummet to random levels, indicating that visual-only cues are insufficient. Furthermore, Figure 4 shows that widely used models such as Qwen2.5-Omni-7B perform closer to random guessing on our benchmark, indicating that OmniVideoBench presents significantly greater challenges than existing multimodal datasets.
@@ 3.1 Baseline Models
We evaluate open-source MLLMs (i.e., Qwen3-Omni series (Xu et al., 2025b), Qwen2.5-Omni series (Xu et al., 2025a), Baichuan-Omni-1.5 (Li et al., 2025), HumanOmni (Zhao et al., 2025), MiniCPM-o (Yao et al., 2024), VideoLLaMA2 (Cheng et al., 2024)), VITA-1.5-7B (Fu et al., 2025), OmniVinci-9B (Ye et al., 2025) and various closed-source MLLMs (i.e., Gemini-2.5-Pro, Gemini-2.5-Flash (Comanici et al., 2025), and Gemini-2.0-Flash). We also evaluate the Qwen2.5-VL series (Bai et al., 2025) and DeepSeek-V3.1 (Liu et al., 2024a).
Models Audio Type Video Duration Avg. Music Sound Speech (0,1] min (1,5] min (5,10] min (10,30] min Omni-Modal Language Models (With Visual and Audio) Gemini-3.0-Pro 52.81 55.17 64.13 62.42 66.18 57.02 59.76 61.79 Gemini-2.5-Pro 38.46 57.72 61.66 57.83 64.43 55.02 55.94 58.90 Gemini-3.0-Flash 49.45 50.34 56.69 58.43 55.10 55.90 52.29 55.10 Gemini-2.5-Flash 39.56 57.04 53.17 55.42 55.10 47.37 52.11 52.40 Gemini-2.0-Flash 29.67 40.27 43.21 49.40 43.15 41.05 34.87 41.50 Qwen3-Omni-30B-A3B 37.36 34.67 39.26 45.78 37.03 38.86 35.11 38.40 OmniVinci-9B 30.77 32.67 32.15 38.55 34.11 30.13 27.10 32.10 Baichuan-Omni-1.5 24.18 31.33 31.36 28.92 31.78 28.38 32.44 30.70 HumanOmni-7B 20.87 31.08 31.61 36.57 29.36 29.60 29.25 30.50 VITA-1.5-7B 25.27 28.57 31.49 31.33 27.41 30.57 33.97 30.50 MiniCPM-o 27.47 28.57 30.24 31.43 28.49 34.53 26.15 29.70 Qwen2.5-Omni-7B 23.07 25.33 30.70 41.57 27.41 25.33 26.72 29.30 VideoLLaMA2-7B 26.37 30.67 29.25 32.00 28.20 29.60 28.29 29.20 Omni-Modal Language Models (Visual Only) Gemini-2.0-Flash 25.27 36.67 30.99 33.73 35.86 32.75 22.48 31.30 Qwen2.5-Omni-7B 27.47 26.67 26.22 28.31 27.11 24.45 25.95 26.40 Visual Language Models (Visual Only) Qwen2.5-VL-32B 32.97 32.00 31.49 38.55 31.20 29.26 30.53 31.80 Qwen2.5-VL-7B 29.67 31.33 29.51 25.90 30.03 31.88 30.15 29.80 Qwen2.5-VL-72B 26.37 29.33 29.91 33.13 30.03 31.88 24.43 29.50 Baseline LLMs DeepSeek-V3.1 28.57 26.17 27.28 30.91 27.57 25.00 26.44 27.60
@@ 3.3 Further Analysis
Models Open-ended QA MCQ
Gemini-2.0-Flash 27.06 41.50
Qwen2.5-Omni-7B 17.25 29.30
Effect of Different Numbers of Frames.
@@ 4 Related Works
Omni-Understanding MLLMs. The development of MLLMs (Chen et al., 2022; Awadalla et al., 2023; Liu et al., 2023; Peng et al., 2025; Yang et al., 2023) began with a foundational focus on integrating the two primary modalities of vision and language. A recent paradigm shift aims to develop Omni-modal MLLMs capable of processing and generating information across an arbitrary combination of modalities (“Any-to-Any"). This approach positions the LLM as a central cognitive engine, unifying diverse data types like audio, video, and text within its semantic space (Liu et al., 2024b; Yuan et al., 2025). This has driven a move from integrating pre-trained unimodal components towards developing “natively multimodal" architectures trained from the ground up, as exemplified by models like GPT-4o (Hurst et al., 2024). This ambition is showcased by state-of-the-art models (Xu et al., 2025a; Zhao et al., 2025; Li et al., 2024b; 2025; Yao et al., 2024; Sun et al., 2025; Liu et al., 2025c; Wu et al., 2025a), which pioneer end-to-end streaming capabilities for simultaneously processing video and audio to generate text and speech. At the forefront of this paradigm, proprietary models like Gemini series (Team, 2024; Comanici et al., 2025) demonstrate pinnacle performance, powered by a natively multimodal design and a massive context window that together unlock superior understanding of complex, interwoven data streams.
MLLM Benchmarks. The landscape of MLLM evaluation has matured significantly, evolving from foundational perception benchmarks (Liu et al., 2024c; Li et al., 2024a; Yu et al., 2024a; b; Chen et al., 2024a; Jiang et al., 2025) to more sophisticated frameworks (He et al., 2025; Du et al., 2025; Wu et al., 2025b). Recent efforts probe deeper cognitive abilities, with MLLM-Bench (Ge et al., 2025) assessing a hierarchy of cognitive skills. MMMU (Yue et al., 2023) and MMMU-Pro (Yue et al., 2024b) challenging models with expert-level, multi-disciplinary reasoning under stricter protocols like vision-only inputs. Simultaneously, evaluation has specialized into high-stakes domains such as finance (Gan et al., 2024) and medicine (Chen et al., 2024b). For video, some benchmarks (Wang et al., 2019; Li et al., 2021; 2023; Fang et al., 2024; Wu et al., 2024) now focus on the critical challenge of long-context temporal understanding (Liu et al., 2025b), revealing key limitations in current models.
@@ 6 Contributions
  • Wentao Wang, Nanjing University